What is the probability of being dealt a royal flush ( A,K,Q,J,10, all the same suit)
Answer #1

number of ways to pick A,K,Q,J,10 from same suit =N(pick one of 4 suits and have A,K,Q,J,10)

=C(4,1) =4

number of ways to have 5 cards from 52 =C(52,5)=2598960

therefore probability of being dealt a royal flush =4/2598960 =1/649740